# Magnon squeezing via reservoir-engineered optomagnomechanics

### v1

@@ I Introduction
The past decade has witnessed the formation and fast development of cavity magnonics, review1 ; review2 ; review3 ; review4 which studies the interaction between microwave cavity photons with collective spin excitations (magnons) in magnetic materials, such as yttrium iron garnet (YIG). One of the most prominent advantages of the magnonic system is that it exhibits an excellent ability to coherently couple with a variety of physical systems, including microwave photons, Huebl2013 ; Tabuchi2014 ; Zhang2014 optical photons, Osada2016 ; Tang2016 ; Haigh2016 vibration phonons, Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 and superconducting qubits, Tabuchi2015 ; DLQ2017 ; DLQ2020 ; Xu2023 etc. The magnonic system therefore has great potential for building a hybrid quantum system with multiple components of different nature, which finds broad applications in quantum information science and technology, review1 ; review2 ; review3 ; review4 quantum sensing, Jing24 and quantum networks. LiPRXQ
@@
Here, we propose to prepare magnonic squeezed states using the novel OMM system. We note that several proposals have been offered to prepare magnonic squeezed states. Specifically, they can be achieved by transferring squeezing from microwaves to magnons, Li2019B two-tone driving of the magnon mode, WangHF2021 ; Qian2024 two-tone driving of a qubit coupled to the magnon mode, Guo2023 and exploiting the anisotropy of the ferromagnet, Kamra2016 ; Sharma2021 the nonlinearity of the magnetostriction, LiNSR and the magnon Kerr effect due to the magnetocrystalline anisotropy, Asjad2023 etc. In this work, we adopt a different approach by applying the idea of reservoir engineering to the OMM system, and consider two scenarios depending on whether the magnomechanical coupling is linear or dispersive. The strong linear coupling is typically the case where the mechanical frequency is close to the magnon frequency and occurs for higher-frequency phonon modes (in the GHz range) associated with thin YIG films, An20 ; Xu21 while the weak dispersive coupling arises for lower-frequency phonon modes (in the MHz range) being much lower than the magnon frequency associated with, e.g., macroscopic YIG spheres. Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 We first generate strong squeezing of the mechanical mode by driving the optical cavity with a two-tone laser field . We then show that in both cases mechanical squeezing can be efficiently transferred to the magnon mode: Stationary magnon squeezing is obtained in the linear coupling case; while transient squeezing is achieved in the dispersive coupling case using a two-step protocol.
@@ II The OMM system with linear magnomechanical interaction
Figure 1: (a) Schematic diagram of the magnon squeezing protocol using the OMM system with a linear magnomechanical coupling. The mechanical mode ($b$) couples to an optical cavity mode ($a$) via the radiation-pressure interaction and to a magnon mode ($m$) via a linear beam-splitter interaction. (b) Mode and drive frequencies used in the protocol. Strong mechanical squeezing is achieved by simultaneously driving the cavity with two laser fields at frequencies $\omega_{\pm}=\omega_{a}\pm\omega_{b}$, and the squeezing is further transferred to the magnon mode via the phonon-magnon linear coupling.
The OMM system consists of optical cavity photons, ferrimagnetic magnons, and vibration phonons induced by magnetostriction, as depicted in Fig. 1(a). Magnons, quanta of spin wave, can couple to phonons via the magnetostrictive interaction. A linear and strong magnon-phonon coupling becomes dominant when the mechanical frequency is high and nearly resonant with the magnon frequency. CMMreview This is typically the case when the magnomechanical sample is a thin YIG film, An20 ; Xu21 and the strong magnon-phonon coupling leads to magnon polarons. Kikkawa The magnomechanical displacement further couples to an optical cavity mode via the radiation pressure or photoelastic effect, COMreview ; Fan2023A which is a dispersive type interaction. Therefore, in the situation of a linear magnon-phonon coupling, the Hamiltonian of the OMM system reads
@@ III Stationary quantum squeezing of mechanical and magnon modes
The physical mechanism of generating quantum squeezing under the two-tone driving field can be explained as follows. Under the RWA, the two driving fields at the blue and red sidebands ($\omega_{\pm}=\omega_{a}\pm\omega_{b}$) of the optical cavity can respectively activate the optomechanical parametric down conversion and beam-splitter interactions, and the effective optomechanical interaction can thus be denoted by $\delta a^{\dagger}(G_{+}\delta b^{\dagger}+G_{-}\delta b)+\mathrm{H.c.}$, cf., Eq. (4). Using the standard squeezing transformation, CMMreview ; Kronwald2013 we introduce the Bogoliubov mode operator for the mechanical mode, $\delta B=\mathrm{cosh}(r)\delta b+\mathrm{sinh}(r)\delta b^{\dagger}$, where the squeezing parameter $r$ is defined as $r=\frac{1}{2}\mathrm{ln}\frac{G_{-}+G_{+}}{G_{-}-G_{+}}$. The effective optomechanical interaction is thus rewritten as $\tilde{G}\delta a^{\dagger}\delta B+\mathrm{H.c.}$, where $\tilde{G}=\sqrt{G_{-}^{2}-G_{+}^{2}}$ denotes the coupling strength between the cavity mode and the Bogoliubov mode. The above cavity-Bogoliubov beam-splitter interaction can lead to the Bogoliubov mode being significantly cooled to its quantum ground state, which corresponds to the squeezed vacuum state of the mechanical mode. By further exploiting the strong phonon-magnon beam-splitter coupling, the squeezing can be efficiently transferred from the mechanical mode to the magnon mode. Consequently, stationary squeezing of the mechanical and magnon modes can be generated under the continuous driving of the two-tone laser field.
@@
where $\langle\delta Q(t)^{2}\rangle$ denotes the variance of the quadrature of one specific mode and $\langle\delta Q(t)^{2}\rangle_{\mathrm{vac}}=\frac{1}{2}$ corresponds to the vacuum fluctuation. In Figs. 2(a) and 2(b), we show the stationary squeezing of the mechanical displacement and magnon phase quadrature, using the following feasible parameters: COMreview ; Carlos2020B $\omega_{m}/2\pi=\omega_{b}/2\pi=10\ \mathrm{GHz}$, cavity resonant wavelength $\lambda=1064\ \mathrm{nm}$, $\kappa_{a}/2\pi=10^{2}\ \mathrm{MHz}$, $\kappa_{m}/2\pi=1\ \mathrm{MHz}$, $\gamma_{b}/2\pi=10^{4}\ \mathrm{Hz}$, $g_{0}/2\pi=10^{2}\ \mathrm{kHz}$, $g_{m}/2\pi=10\ \mathrm{MHz}$, $T=10\ \mathrm{mK}$, and a fixed $G_{-}/2\pi=15\ \mathrm{MHz}$, corresponding to a pump power of $P_{-}\simeq 26.41\ \mathrm{mW}$ for the driving field at frequency $\omega_{-}=\omega_{a}-\omega_{b}$. For the stability reason, we consider $G_{+}<G_{-}$. Qian2024 Clearly, there exists an optimal ratio of $G_{+}/G_{-}$ for the maximal mechanical (magnon) squeezing, as a result of the trade-off between the following two effects: On the one hand, a sufficiently large difference between the two couplings $G_{+}$ and $G_{-}$ is required to have a sufficiently large cooling rate $\tilde{G}$ of the Bogoliubov mode; On the other hand, almost equal couplings $G_{+}\approx G_{-}$ is needed to obtain an extremely large squeezing of the mechanical mode. Figures 2(c) and 2(d) illustrate the corresponding Wigner function of the squeezed mechanical and magnon modes, with an optimal ratio of $G_{+}/G_{-}$.
Figure 3: The degree of stationary magnon squeezing $S_{m}$ versus $G_{+}/G_{-}$, at different bath temperatures of $T=0.01$ K (solid), $T=0.5$ K (dashed), and $T=1.0$ K (dot-dahed). The other parameters are the same as in Fig. 2(b).
In Fig. 3, we study the degree of squeezing of the magnon mode $S_{m}$ versus the ratio $G_{+}/G_{-}$ at different bath temperatures. As the temperature rises, the system suffers more thermal noises and a larger cooling rate $\tilde{G}$ is thus needed to efficiently cool the Bogoliubov mode to its ground state, which implies a larger difference between $G_{+}$ and $G_{-}$, i.e., a decreasing optimal ratio $G_{+}/G_{-}$ (for a fixed $G_{-}$). This, in turn, reduces the degree of squeezing that can be achieved in the steady state. Nevertheless, we see that stationary magnon squeezing can be achieved for bath temperature above 1 K, manifesting the robustness of the protocol against thermal noises.
@@ IV Magnon squeezing with dispersive magnomechanical interaction
The above two sections study stationary magnonic squeezed states based on the OMM system with a linear magnomechanical coupling. However, for lower-frequency mechanical modes, e.g., of a sub-millimeter sized YIG sphere Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 ; ShenZ2022 or a micron sized YIG bridge, Heyroth2019 ; Fan2023A ; Kansanen2021 the magnomechanical interaction is dominated by the nonlinear dispersive coupling. CMMreview In this case, the OMM system has both nonlinear opto- and magnomechanical interactions. The latter weak dispersive coupling hinders an efficient phonon-to-magnon squeezing transfer (which prefers a linear strong coupling as discussed in Secs. II and III). Nevertheless, we show in this section that transient moderate squeezing of the magnon mode can still be achieved by using a two-step protocol, as depicted in Fig. 4(a).
Figure 4: (a) Schematic diagram of the two-step magnon squeezing protocol using the OMM system with a dispersive magnomechanical coupling. (b) Mode and drive frequencies used in the two-step protocol. In the first step, the optical cavity is simultaneously driven by two laser fields at frequencies $\omega_{\pm}=\omega_{a}\pm\omega_{b}$ to generate stationary squeezing of the mechanical mode. In the second step, the squeezing is transferred from the mechanical mode to the magnon mode via the magnon-phonon state-swap interaction, realized by driving the magnon mode with a red-detuned microwave field.
@@
Figure 5: (a) The degree of squeezing $S_{b}$ of the mechanical mode versus $G_{+}/G_{-}$. (b) The Wigner function of the squeezed mechanical mode, with $G_{+}/G_{-}=0.95$. See text for the other parameters.
In Fig. 5(a), we plot the degree of squeezing $S_{b}$ of the mechanical mode versus $G_{+}/G_{-}$, using experimentally feasible parameters: $\omega_{b}/2\pi=10^{2}\ \mathrm{MHz}$, cavity resonant wavelength $\lambda=1064\ \mathrm{nm}$, $\kappa_{a}/2\pi=2\ \mathrm{MHz}$, $\gamma_{b}/2\pi=10^{2}\ \mathrm{Hz}$, $g_{0}/2\pi=1\ \mathrm{kHz}$, and $T=10\ \mathrm{mK}$. We take $G_{-}/2\pi=0.3\ \mathrm{MHz}$ corresponding to a power $P_{-}\simeq 0.53$ mW of the driving field at frequency $\omega_{-}=\omega_{a}-\omega_{b}$. A strong squeezing (13 dB) of the mechanical displacement is achieved at the optimal ratio $G_{+}/G_{-}\simeq 0.95$, corresponding to a power $P_{+}\simeq 0.48$ mW of the driving field at frequency $\omega_{+}=\omega_{a}+\omega_{b}$. Figure 5(b) shows the corresponding Wigner function of the squeezed mechanical mode.

### v2

@@ I Introduction
The past decade has witnessed the formation and fast development of cavity magnonics, review1 ; review2 ; review3 ; review4 which studies the interaction between microwave cavity photons with collective spin excitations (magnons) in magnetic materials, such as yttrium iron garnet (YIG). One of the most prominent advantages of the magnonic system is that it exhibits an excellent ability to coherently couple with a variety of physical systems, including microwave photons, Huebl2013 ; Tabuchi2014 ; Zhang2014 optical photons, Osada2016 ; Tang2016 ; Haigh2016 vibration phonons, Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 and superconducting qubits, Tabuchi2015 ; DLQ2017 ; DLQ2020 ; Xu2023 etc. The magnonic system therefore has great potential for building a hybrid quantum system with multiple components of different nature, which finds broad applications in quantum information science and technology, review1 ; review2 ; review3 ; review4 quantum sensing, Jing2024 and quantum networks. LiPRXQ
@@
Here, we propose to prepare magnonic squeezed states using the novel OMM system. We note that several proposals have been offered to prepare magnonic squeezed states. Specifically, they can be achieved by transferring squeezing from microwaves to magnons, Li2019B two-tone driving of the magnon mode, WangHF2021 ; Qian2024 two-tone driving of a qubit coupled to the magnon mode, Guo2023 and exploiting the anisotropy of the ferromagnet, Kamra2016 ; Sharma2021 the nonlinearity of the magnetostriction, LiNSR and the magnon Kerr effect due to the magnetocrystalline anisotropy, Asjad2023 etc. In this work, we adopt a different approach by applying the idea of reservoir engineering to the OMM system, and consider two scenarios depending on whether the magnomechanical coupling is linear or dispersive. The strong linear coupling is typically the case where the mechanical frequency is close to the magnon frequency and occurs for higher-frequency phonon modes (in the GHz range) associated with thin YIG films, An20 ; Xu21 while the weak dispersive coupling arises for lower-frequency phonon modes (in the MHz range) being much lower than the magnon frequency associated with, e.g., macroscopic YIG spheres. Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 We first generate strong squeezing of the mechanical mode by driving the optical cavity with a two-tone laser field . We then show that in both cases mechanical squeezing can be efficiently transferred to the magnon mode: Stationary magnon squeezing is obtained in the linear coupling case; while transient squeezing is achieved in the dispersive coupling case using a two-step protocol. The quantum squeezing finds many important applications in quantum information processing and quantum sensing in the fields of cavity magnomechanics Sharma2021 ; Sun2021 ; Lu2023 ; Wan2024 ; Jing2024 ; ZhangQF2024 ; Qian2024 and optomechanics. He2013 ; Qiu2020 ; Lu2024 ; Tan2020 ; Jiao2024 ; LIGO2013 ; Xu2014 ; Peano2015 ; Motazedifard2016 ; Clark2016 ; ZhaoW2020
@@ II The OMM system with linear magnomechanical interaction
Figure 1: (a) Schematic diagram of the magnon squeezing protocol using the OMM system with a linear magnomechanical coupling. The mechanical mode ($b$) couples to an optical cavity mode ($a$) via the radiation-pressure interaction and to a magnon mode ($m$) via a linear beam-splitter interaction. Below are the typical optomechanical system and the magnomechanical system, e.g., a YIG film, with a linear coupling. (b) Mode and drive frequencies used in the protocol. Strong mechanical squeezing is achieved by simultaneously driving the cavity with two laser fields at frequencies $\omega_{\pm}=\omega_{a}\pm\omega_{b}$, and the squeezing is further transferred to the magnon mode via the phonon-magnon linear coupling.
The OMM system consists of optical cavity photons, ferrimagnetic magnons, and vibration phonons induced by magnetostriction, as depicted in Fig. 1(a). Magnons, as quanta of spin wave, are collective excitations of a large number of spins in a ferrimagnet, such as YIG. Originating from the magnetoelastic energy of the magnetic material, magnons can couple to phonons via the magnetostrictive interaction, which describes the interaction between magnetization and elastic strain of the magnetic material. CMMreview A linear and strong beam-splitter (BS) type magnon-phonon coupling becomes dominant when the mechanical frequency is high and close to the magnon frequency in the GHz range. This is typically the case when the magnomechanical sample is a thin YIG film, An20 ; Xu21 as depicted in Fig. 1(a), and the strong magnon-phonon coupling leads to magnon polarons. Kikkawa The magnomechanical displacement further couples to an optical cavity mode via the radiation pressure or photoelastic effect, COMreview ; Fan2023A which can affect the resonance frequency of the cavity field, manifesting a dispersive type interaction. Therefore, in the situation of a linear magnon-phonon coupling, the Hamiltonian of the OMM system reads
@@ III Stationary quantum squeezing of mechanical and magnon modes
The physical mechanism of generating quantum squeezing under the two-tone driving field can be explained as follows. Under the RWA, the two driving fields at the blue and red sidebands ($\omega_{\pm}=\omega_{a}\pm\omega_{b}$) of the optical cavity can respectively activate the optomechanical parametric down conversion and beam-splitter interactions, and the effective optomechanical interaction can thus be denoted by $\delta a^{\dagger}(G_{+}\delta b^{\dagger}+G_{-}\delta b)+\mathrm{H.c.}$, cf., Eq. (4). Using the standard squeezing transformation, CMMreview ; Kronwald2013 we introduce the Bogoliubov mode operator for the mechanical mode, $\delta B=\mathrm{cosh}(r)\delta b+\mathrm{sinh}(r)\delta b^{\dagger}$, where the squeezing parameter $r$ is defined as $r=\frac{1}{2}\mathrm{ln}\frac{G_{-}+G_{+}}{G_{-}-G_{+}}$. The effective optomechanical interaction is thus rewritten as $\tilde{G}\delta a^{\dagger}\delta B+\mathrm{H.c.}$, where $\tilde{G}=\sqrt{G_{-}^{2}-G_{+}^{2}}$ denotes the coupling strength between the cavity mode and the Bogoliubov mode. The above cavity-Bogoliubov beam-splitter interaction can lead to the Bogoliubov mode being significantly cooled to its quantum ground state, which corresponds to the squeezed state of the mechanical mode. By further exploiting the strong phonon-magnon beam-splitter coupling, the squeezing can be efficiently transferred from the mechanical mode to the magnon mode. Consequently, stationary squeezing of the mechanical and magnon modes can be generated under the continuous driving of the two-tone laser field.
@@
where $\langle\delta Q(t)^{2}\rangle$ denotes the variance of the quadrature of one specific mode and $\langle\delta Q(t)^{2}\rangle_{\mathrm{vac}}=\frac{1}{2}$ corresponds to the vacuum fluctuation. In Figs. 2(a) and 2(b), we show the stationary squeezing of the mechanical displacement and magnon phase quadrature, using the following feasible parameters: An20 ; Xu21 ; Chan2011 ; Chan2012 ; Hong2017 $\omega_{m}/2\pi=\omega_{b}/2\pi=10\ \mathrm{GHz}$, cavity resonant wavelength $\lambda=1064\ \mathrm{nm}$, $\kappa_{a}/2\pi=10^{2}\ \mathrm{MHz}$, $\kappa_{m}/2\pi=1\ \mathrm{MHz}$, $\gamma_{b}/2\pi=10^{4}\ \mathrm{Hz}$, $g_{0}/2\pi=10^{2}\ \mathrm{kHz}$, $g_{m}/2\pi=10\ \mathrm{MHz}$, $T=10\ \mathrm{mK}$, and a fixed $G_{-}/2\pi=15\ \mathrm{MHz}$, corresponding to a pump power of $P_{-}\simeq 26.41\ \mathrm{mW}$ for the driving field at frequency $\omega_{-}=\omega_{a}-\omega_{b}$. For the stability reason, we consider $G_{+}<G_{-}$. Qian2024 Clearly, there exists an optimal ratio of $G_{+}/G_{-}$ for the maximal mechanical (magnon) squeezing, as a result of the trade-off between the following two effects: On the one hand, a sufficiently large difference between the two couplings $G_{+}$ and $G_{-}$ is required to have a sufficiently large cooling rate $\tilde{G}$ of the Bogoliubov mode; On the other hand, almost equal couplings $G_{+}\approx G_{-}$ is needed to obtain an extremely large squeezing of the mechanical mode. Figures 2(c) and 2(d) illustrate the corresponding Wigner function of the squeezed mechanical and magnon modes, with an optimal ratio of $G_{+}/G_{-}$.
Figure 3: (a) The degree of stationary magnon squeezing $S_{m}$ and (b) effective phonon number of the Bogoliubov mode $N_{B}$ versus $G_{+}/G_{-}$, at different bath temperatures of $T=0.01$ K (solid), $T=0.5$ K (dashed), and $T=1.0$ K (dot-dahed). The other parameters are the same as in Fig. 2(b).
In Fig. 3(a), we study the degree of squeezing of the magnon mode $S_{m}$ versus the ratio $G_{+}/G_{-}$ at different bath temperatures. As the temperature rises, the system suffers more thermal noises and a larger cooling rate $\tilde{G}$ is thus needed to efficiently cool the Bogoliubov mode to its ground state, which implies a larger difference between $G_{+}$ and $G_{-}$, i.e., a decreasing optimal ratio $G_{+}/G_{-}$ (for a fixed $G_{-}$). This, in turn, reduces the degree of squeezing that can be achieved in the steady state. Nevertheless, we see that stationary magnon squeezing can be achieved for bath temperature above 1 K, manifesting the robustness of the protocol against thermal noises. In Fig. 3(b), we also plot the effective phonon number of the Bogoliubov mode $N_{B}$ versus the ratio $G_{+}/G_{-}$ at different bath temperatures. At the optimal ratio $G_{+}/G_{-}$ for each temperature, we obtain $N_{B}\simeq 0.30$ for $T=0.01$ K; 0.39 for $T=0.5$ K; and 0.50 for $T=1.0$ K, indicating that the Bogoliubov mode is cooled (close) to its quantum ground state in all the situations. Note that for $G_{+}/G_{-}$ lower than the optimal ratio, the Bogoliubov mode is also in the ground state, but due to a smaller $G_{+}$ ($G_{-}$ is fixed), which leads to a smaller squeezing parameter $r$, the degree of squeezing of the mechanical mode (and thus the magnon mode) reduces.
@@ IV Magnon squeezing with dispersive magnomechanical interaction
The above two sections study stationary magnonic squeezed states based on the OMM system with a linear magnomechanical coupling. However, for lower-frequency mechanical modes, e.g., of a sub-millimeter sized YIG sphere Zhang2016 ; Li2018 ; Potts2021 ; Shen2022 ; ShenZ2022 or a micron sized YIG bridge, Heyroth2019 ; Fan2023A ; Kansanen2021 ; Arisawa2019 the magnomechanical interaction is dominated by the nonlinear dispersive coupling. CMMreview In this case, the magnetostrictive interaction gives rise to a mechanical displacement of the YIG sample (see Fig. 4(a)), which is proportional to the magnon excitation number. The OMM system then has both nonlinear opto- and magnomechanical interactions. The latter weak dispersive coupling hinders an efficient phonon-to-magnon squeezing transfer (which prefers a linear strong coupling as discussed in Secs. II and III). Nevertheless, we show in this section that transient moderate squeezing of the magnon mode can still be achieved by using a two-step protocol, as depicted in Fig. 4(a).
Figure 4: (a) Schematic diagram of the two-step magnon squeezing protocol using the OMM system with a dispersive magnomechanical coupling. Below are the typical optomechanical system and the magnomechanical system, e.g., a YIG micro bridge, with a dispersive coupling. (b) Mode and drive frequencies used in the two-step protocol. In the first step, the optical cavity is simultaneously driven by two laser fields at frequencies $\omega_{\pm}=\omega_{a}\pm\omega_{b}$ to generate stationary squeezing of the mechanical mode. In the second step, the squeezing is transferred from the mechanical mode to the magnon mode via the magnon-phonon state-swap interaction, realized by driving the magnon mode with a red-detuned microwave field.
@@
Figure 5: (a) The degree of squeezing $S_{b}$ of the mechanical mode versus $G_{+}/G_{-}$. (b) The Wigner function of the squeezed mechanical mode, with $G_{+}/G_{-}=0.95$. (c) The effective phonon number of the Bogoliubov mode versus $G_{+}/G_{-}$. See text for the other parameters.
In Fig. 5(a), we plot the degree of squeezing $S_{b}$ of the mechanical mode versus $G_{+}/G_{-}$, using experimentally feasible parameters: Arcizet2006 ; Favero2007 ; Simon2009b $\omega_{b}/2\pi=10^{2}\ \mathrm{MHz}$, cavity resonant wavelength $\lambda=1064\ \mathrm{nm}$, $\kappa_{a}/2\pi=2\ \mathrm{MHz}$, $\gamma_{b}/2\pi=10^{2}\ \mathrm{Hz}$, $g_{0}/2\pi=1\ \mathrm{kHz}$, and $T=10\ \mathrm{mK}$. We take $G_{-}/2\pi=0.3\ \mathrm{MHz}$ corresponding to a power $P_{-}\simeq 0.53$ mW of the driving field at frequency $\omega_{-}=\omega_{a}-\omega_{b}$. A strong squeezing (13 dB) of the mechanical displacement is achieved at the optimal ratio $G_{+}/G_{-}\simeq 0.95$, corresponding to a power $P_{+}\simeq 0.48$ mW of the driving field at frequency $\omega_{+}=\omega_{a}+\omega_{b}$. Figure 5(b) shows the corresponding Wigner function of the squeezed mechanical mode. Similarly, we plot the effective phonon number of the Bogoliubov mode versus $G_{+}/G_{-}$ in Fig. 5(c). The Bogoliubov mode is also cooled to the ground state with $N_{B}\simeq 0.23$ at the optimal ratio $G_{+}/G_{-}\simeq 0.95$.